# Counterfactual Visual Explanation via Causally-Guided Adversarial Steering

### v1

@@ Abstract
Recent work on counterfactual visual explanations has contributed to making artificial intelligence models more explainable by providing visual perturbation to flip the prediction. However, these approaches neglect the causal relationships and the spurious correlations behind the image generation process, which often leads to unintended alterations in the counterfactual images and renders the explanations with limited quality. To address this challenge, we introduce a novel framework CECAS, which first leverages a causally-guided adversarial method to generate counterfactual explanations. It innovatively integrates a causal perspective to avoid unwanted perturbations on spurious factors in the counterfactuals. Extensive experiments demonstrate that our method outperforms existing state-of-the-art approaches across multiple benchmark datasets and ultimately achieves a balanced trade-off among various aspects of validity, sparsity, proximity, and realism.
@@ 1 Introduction
Despite the impressive performance of deep neural networks (DNNs) across domains such as language, vision, and control [1, 2, 3], explaining their predictions remains challenging due to their non-linear, overparameterized nature [4]. This challenge is particularly critical in safety-sensitive applications like healthcare, finance, and autonomous driving [5]. As a result, explainable AI (XAI) has gained momentum [6], with counterfactual explanations (CEs) [7] emerging as a powerful approach. CEs aim to answer: What minimal change to an input would alter the model’s prediction? They are valued for producing intuitive and preferably actionable insights that can support decision auditing and planning. As a powerful XAI tool [7], CE has been particularly flourishing in computer vision [8, 9, 10, 11]. Unlike attribution-based methods that highlight pixels associated with the output, counterfactuals in images simulate realistic alternative scenarios, allowing us to ask: What would the image need to look like for the model to decide differently? For example, modifying a few facial features to change a classification from “neutral” to “happy” provides insight into the model’s decision boundary. Successful visual counterfactuals not only satisfy conventional CE criteria such as being valid (flip the prediction), proximal (close to the original image), and sparse (modify as few features as possible) [12, 13], but also are semantically meaningful and perceptually natural, avoiding unrealistic artifacts. These methods offer intuitive and human-understandable explanations, making them especially valuable for analyzing high-dimensional visual models.
@@
While counterfactual visual explanation has seen rapid progress in recent years [8, 9, 10, 11], several critical aspects of the problem remain overlooked. A key limitation lies in the lack of a principled understanding of the underlying causal mechanisms in the data, as well as the true target of explanation. Most existing methods [14, 15, 16] optimize directly over the entire input feature space in search of perturbations, which often leads to changes in spurious factors that are correlated with the label but not causally related. While such perturbations may achieve a prediction flip, they rarely reflect meaningful or actionable changes in the real world and can even produce misleading or harmful outcomes. For example, in autonomous driving scenarios, changing the shape of a traffic sign (spurious factor) to alter a “stop” vs. “go” prediction is far less informative than changing the content of the sign, which is causally relevant to the model’s decision. More importantly, conventional CE strategies often result in low-quality and unrealistic images. The pixel space of images is vast, and unconstrained perturbations can introduce noise or artifacts that are visually implausible or semantically incoherent. Recent advances [9, 11] have begun to address this by leveraging generative models, particularly diffusion-based methods, which operate in semantically meaningful latent spaces and are better equipped to produce realistic counterfactuals. These approaches benefit from the ability of diffusion models to suppress high-frequency noise and structure perturbations more coherently. However, most current methods still rely heavily on predefined loss functions, such as $l_{1}$ loss constraining the difference between original and counterfactual images, to enforce sparsity and proximity. While effective at a global level, these constraints often lead to unwanted changes in local artifacts such as color shifts, shape distortions, or the emergence of spurious attributes [17]. This results in a persistent trade-off between validity, proximity, sparsity, and realism, making it difficult to generate counterfactuals that are both high-quality and causally meaningful. To the best of our knowledge, no existing method well addresses these challenges, underscoring the need for a new approach that incorporates causal reasoning into the generation of counterfactual visual explanations.
@@
To address the aforementioned challenges, we propose a novel framework, Counterfactual Explanation via Causally Adversarial Steering (CECAS), which leverages a causality-guided adversarial method to generate counterfactual explanations that are more semantically faithful and causally grounded. Adversarial attacks have been an effective tool to generate CEs due to their common aim to minimally alter an input to change a model’s prediction [18]. Uniquely, our approach integrates causality into the adversarial objective: While the adversarial component ensures prediction flips (validity), the causal guidance directs perturbations toward features that are causally relevant to the label. This promotes sparsity, proximity, and realism, by avoiding changes to spurious or irrelevant features. As a result, our method produces high-quality, semantically meaningful counterfactual images that align with real-world interventions. We conclude our contributions as follows:
@@ 2 Related Work
Explainable AI (XAI) XAI refers to a set of techniques aimed at enhancing the transparency and explainability of AI models—making their decisions and internal mechanisms more understandable to humans [5]. The field is commonly divided into two major categories: ad-hoc and post-hoc methods. The key difference between them lies in the stage at which they are implemented [19]. The former is often deployed during the training phase and advocates the design of inherently interpretable architectures [20, 21, 22, 23, 24]. In contrast, the latter focuses on understanding and explaining the trained models without modifying their internal structure [25, 26, 27]. The post-hoc category includes global and local explanations, distinguished by the scope of the explanations they provide [5]. Global explanation provides a comprehensive description of the model [28], while the latter focuses on explaining the model’s prediction for a specific instance [29]. The studied problem of our paper belongs to local explanations. This line of work mainly focuses on the following three directions: saliency map [30, 31, 32], concept attribution [33, 34, 35], and model distillation [36, 37].
@@
Counterfactual Explanations As a branch of post-hoc explanations, CE has emerged as a highly influential approach in recent years [18], especially in computer vision [11]. Some methods generate counterfactual samples by locating key regions that distinguish two contrasting images and substituting the relevant region from one image into the other [38, 39]. Other methods utilize the gradient of the input image with respect to the target label while also leveraging robust models to generate human-recognizable modifications [40, 41]. Additional approaches, including prototype-based algorithms [42] and invertible convolutional networks [43], have also shown promising abilities in generating counterfactual examples. Generative techniques are inherently well-suited for CE. With their rapid development, an increasing number of related works have emerged, which can be broadly categorized into conditional [44, 45] and unconditional approaches [46, 10]. Diffusion models (DDPM) are cutting-edge generative models. Jeanneret et al. [9] and [11] both utilize DDPM to generate counterfactual images. The difference lies in that the former modifies the sampling process of the DDPM, while the latter employs DDPM as a regularizer.
@@
Adversarial Attacks Adversarial attacks and CEs share a similar objective: flipping the model’s prediction, which has led some studies to focus on exploring the relationship between them [47, 48, 18]. However, the key difference lies in the fact that adversarial attacks on undefended models often result in imperceptible noise rather than semantically meaningful changes. Adversarial attacks include white-box and black-box settings. In white-box attacks, the attacker has access to the model’s architecture and data [49]. It typically generates adversarial examples by utilizing the gradient of the loss function with respect to the input [50, 51, 52]. In contrast, black-box attacks [53, 54] restrict the attacker to accessing only samples from an oracle or pairs of inputs and corresponding outputs.
@@
Adversarial Defense The development of adversarial attacks has, in turn, stimulated significant progress in adversarial defense. Recent works can be primarily divided into two categories: certified defense and empirical defense. Certified defense [55, 56, 57] offer provable robustness by ensuring that the model’s predictions remain stable under perturbations constrained within a specified norm ball. However, empirical defense [50, 51, 58], particularly adversarial training [51], is still highly effective. It typically generates adversarial samples by using an adversarial attack first and then incorporates such samples into the training process. It is worth noting that recent work has begun to integrate causality into adversarial defense strategies [59, 60, 61, 62], sharing a common objective with our work.
@@ 3 Preliminaries
Causal View on Image Generation Figure 1 is a well-recognized causal graph $\mathcal{G}$ [62] that can represent the generation processes of both the natural and counterfactual images. As shown, $C$ denotes the causal factors that directly determine the label $Y$ and predicted label $\hat{Y}$. $S$ denotes the spurious factors, which are independent of $C$, as reflected in the $v$-structure $C\rightarrow{X}\leftarrow{S}$. $X$ is the natural image, and $\Delta$ represents the semantically meaningful perturbation; together, they form the counterfactual image $X^{\prime}$. $\theta$ is the set of parameters of the DNN. It is evident from the graph that the perturbation $\Delta$ is determined by $\hat{Y}$, $X$ and $\theta$. The inclusion of $\theta$ is due to the use of white-box attacks, such as projected gradient descent (PGD). Note that $S$ can also have an influence on $Y$ because of the spurious paths in $\mathcal{G}$. When $X$ is given, there is a spurious path ${Y}\leftarrow{C}\rightarrow{X}\leftarrow{S}$ associate $Y$ and $S$.
Denoising Diffusion Probabilistic Models DDPM [63] can filter out the noise introduced by traditional adversarial attacks without changing the classifier’s weights. It relies on two Markov chains, corresponding to the forward and reverse processes, respectively. The forward process gradually adds Gaussian noise to the data at the state of $t$ to generate data at the next state $t+1$. In practice, the data at time step $t$, denoted as $x_{t}$ can be directly obtained from the clean data $x_{0}$ through
@@ 4 Methodology
We propose the CECAS framework, which consists of two key components, as shown in Figure 2. The first component is the initial generation of CE images, achieved through a novel causally guided adversarial method. The second component is the causal refinement of the generated counterfactuals, where we adopt an image inpainting approach to better preserve the details of the unmodified regions of the original image. We detail them in the following sections.
@@ 4.1 Initial Causally-Guided Counterfactual Visual Explanation Generation
The first and central component of the proposed framework is a causally guided adversarial method to generate counterfactual visual explanations. As previously mentioned, adversarial attacks and CE share a common goal: altering a model’s prediction through minimal input perturbations. Inspired by prior work bridging the two domain [18, 9, 64], we adopt adversarial strategies—such as PGD attacks—to generate image-based counterfactuals. The procedure is outlined as follows:
@@
However, relying solely on the above adversarial attack-based method is often overly idealized, as these methods typically impose no constraints on spurious factors. As discussed in the Introduction, although the explained model may rely on such factors during prediction, perturbing them often does not yield meaningful or actionable explanations. Furthermore, altering these spurious factors can introduce artifacts or noise, which severely compromises the core goals of CE of providing realistic and constructive explanations.
To address this issue, it is essential to constrain changes to spurious factors. Given the central role of gradients in the above CE generation process, we aim to identify a new gradient direction—one that enables effective modification of the image while preserving spurious factors unchanged. Since our goal is to explain the classifier $f$, we are not allowed to make any modifications to it. Considering the architecture of the explained model $f$, a DNN commonly viewed—following [65]— as a combination of a feature extraction module and a linear classifier module, we learn to capture the spurious factors out of the representations $R$ learned by the feature extraction module (details of how to capture $S$ from $R$ are given in the final part of this subsection). With the captured $S$, we use the original data to train an auxiliary classifier $g$ which predicts the label $Y$ based on the spurious factor. More specifically, after training, $g$ can make predictions as follows: $\hat{z}=g(s),\hat{z}^{\prime}=g(s^{\prime})$, where $\hat{z}$ denotes the logit predicted by $g$ from the spurious factor $s$ in the original image, while $\hat{z}^{\prime}$ is the logit predicted by $g$ from the spurious factor $s^{\prime}$ in the modified image. We can consider $z$ to capture the spurious factors that are most relevant to the label $Y$—that is, the spurious features most likely to be unintentionally altered during counterfactual generation. On top of this, we introduce a penalty loss $L_{S}$ that encourages these spurious factors to remain invariant during the counterfactual generation process, and thereby we compute the gradient of the penalty loss with respect to the image and incorporate its direction into the guidance for image modification:
@@
where $L_{S}$ can typically be implemented as the soft cross-entropy loss, which takes two logits as input. The primary purpose of this loss is to align the spurious distributions between the natural and counterfactual representations by penalizing the deviation of spurious factors in the counterfactuals from those in the original inputs. In summary, the generation of counterfactual images involves modifying the images via adversarial attack, while preserving the spurious factors through a causal perspective.
Furthermore, inspired by the effectiveness of diffusion models in enhancing CE image quality [9, 64], we integrate diffusion models to filter out high-frequency, semantically meaningless noise introduced by adversarial attacks. Specifically, we apply the forward process of DDPM to add Gaussian noise to the image first, and then use the reverse process to denoise it before feeding it into the classifier for loss computation. In this work, we define the joint classifier $F$ as the combination of DDPM and the base classifier $f$. The generation of counterfactual images can be simplified and reformulated in an optimization form:
@@
The only remaining problem is how to obtain the spurious factor $S$. Fortunately, numerous recent works have explored disentangling causal factors from spurious ones [66, 62, 67]. Building on the causal model illustrated in Figure 1, we adopt the widely used assumption that the causal factor $C$ is independent of the spurious factor $S$. Leveraging this independence, we employ a subspace projection technique to project $C$ and $S$ to orthogonal subspaces and thereby extract the spurious factor $S$, as proposed in Zhang et al. [62], this technique is supported by solid theoretical foundations. Notably, our framework is not limited to this specific implementation and can be extended or replaced by other effective strategies for extracting $S$.
4.2 Causally-Guided Post Refinement

### v2

@@ Abstract
Recent work on counterfactual visual explanations has contributed to making artificial intelligence models more explainable by providing visual perturbations to flip the prediction. However, these approaches neglect the causal relationships and the spurious correlations behind the image generation process, which often leads to unintended alterations in the counterfactual images and renders the explanations with limited quality. To address this challenge, we introduce a novel framework CECAS, which leverages a causally-guided adversarial method to generate counterfactual explanations. It innovatively integrates a causal perspective to avoid unwanted perturbations on spurious factors in the counterfactuals. Extensive experiments demonstrate that our method outperforms existing state-of-the-art approaches across multiple benchmark datasets and ultimately achieves a balanced trade-off among various aspects of validity, sparsity, proximity, and realism.
@@ 1 Introduction
Despite the impressive performance of deep neural networks (DNNs) across domains such as language, vision, and control (Devlin et al., 2019; Bojarski et al., 2016; Silver et al., 2017), explaining their predictions remains challenging due to their non-linear, overparameterized nature (Li et al., 2022). This challenge is particularly critical in safety-sensitive applications like healthcare, finance, and autonomous driving (Mersha et al., 2024). As a result, explainable AI (XAI) has gained momentum (Dwivedi et al., 2023), with counterfactual explanations (CEs) Wachter et al. (2017) emerging as a powerful approach. CEs aim to answer: What minimal change to an input would alter the model’s prediction? They are valued for producing intuitive and preferably actionable insights that can support decision auditing and planning. As a powerful XAI tool (Wachter et al., 2017), CE has been particularly flourishing in computer vision (Jacob et al., 2022; Jeanneret et al., 2022; Rodriguez et al., 2021; Jeanneret et al., 2023). Unlike attribution-based methods that highlight pixels associated with the output, counterfactuals in images simulate realistic alternative scenarios, allowing us to ask: What would the image need to look like for the model to decide differently? For example, modifying a few facial features to change a classification from “neutral” to “happy” provides insight into the model’s decision boundary. Successful visual counterfactuals not only satisfy conventional CE criteria such as being valid (flip the prediction), proximal (close to the original image), and sparse (modify as few features as possible) (Mothilal et al., 2020; Russell, 2019), but also are semantically meaningful and perceptually natural, avoiding unrealistic artifacts. These methods offer intuitive and human-understandable explanations, making them especially valuable for analyzing high-dimensional visual models.
@@
While counterfactual visual explanation has seen rapid progress in recent years Jacob et al. (2022); Jeanneret et al. (2022); Rodriguez et al. (2021); Jeanneret et al. (2023), several critical aspects of the problem remain overlooked. A key limitation lies in the lack of a principled understanding of the underlying causal mechanisms in the data, as well as the true target of explanation. Most existing methods (Petsiuk et al., 2018; 2021; Vasu & Long, 2020) optimize directly over the entire input feature space in search of perturbations, which often leads to changes in spurious factors that are correlated with the label but not causally related. While such perturbations may achieve a prediction flip, they rarely reflect meaningful or actionable changes in the real world and can even produce misleading or harmful outcomes. For example, in autonomous driving scenarios, changing the shape of a traffic sign (spurious factor) to alter a “stop” vs. “go” prediction is far less informative than changing the content of the sign, which is causally relevant to the model’s decision. More importantly, conventional CE strategies often result in low-quality and unrealistic images. The pixel space of images is vast, and unconstrained perturbations can introduce noise or artifacts that are visually implausible or semantically incoherent. Recent advances Jeanneret et al. (2022; 2023) have begun to address this by leveraging generative models, particularly diffusion-based methods, which operate in semantically meaningful latent spaces and are better equipped to produce realistic counterfactuals. These approaches benefit from the ability of diffusion models to suppress high-frequency noise and structure perturbations more coherently. However, most current methods still rely heavily on predefined loss functions, such as $l_{1}$ loss constraining the difference between original and counterfactual images, to enforce sparsity and proximity. While effective at a global level, these constraints often lead to unwanted changes in local artifacts such as color shifts, shape distortions, or the emergence of spurious attributes (Andonian et al., 2021). This results in a persistent trade-off between validity, proximity, sparsity, and realism, making it difficult to generate counterfactuals that are both high-quality and causally meaningful. To the best of our knowledge, no existing method well addresses these challenges, underscoring the need for a new approach that incorporates causal reasoning into the generation of counterfactual visual explanations.
@@
To address the aforementioned challenges, we propose a novel framework, Counterfactual Explanation via Causally Adversarial Steering (CECAS), which leverages a causality-guided adversarial method to generate counterfactual explanations that are more semantically faithful and causally grounded. Adversarial attacks have been an effective tool to generate CEs due to their common aim to minimally alter an input to change a model’s prediction Pawelczyk et al. (2022). Uniquely, our approach integrates causality into the adversarial objective: While the adversarial component ensures prediction flips (validity), the causal guidance directs perturbations toward features that are causally relevant to the label. This promotes sparsity, proximity, and realism, by avoiding changes to spurious or irrelevant features. As a result, our method produces high-quality, semantically meaningful counterfactual images that align with real-world interventions. We conclude our contributions as follows:
@@ 2 Related Work
Explainable AI (XAI) XAI refers to a set of techniques aimed at enhancing the transparency and explainability of AI models—making their decisions and internal mechanisms more understandable to humans (Mersha et al., 2024). The field is commonly divided into two major categories: ad-hoc and post-hoc methods. The key difference between them lies in the stage at which they are implemented (Guidotti et al., 2018). The former is often deployed during the training phase and advocates the design of inherently interpretable architectures (Alaniz et al., 2021; Alvarez Melis & Jaakkola, 2018; Bohle et al., 2021; Huang & Li, 2020; Nauta et al., 2021). In contrast, the latter focuses on understanding and explaining the trained models without modifying their internal structure (Selvaraju et al., 2017; Shrikumar et al., 2017; Simonyan et al., 2013). The post-hoc category includes global and local explanations, distinguished by the scope of the explanations they provide (Mersha et al., 2024). Global explanation provides a comprehensive description of the model (Lipton, 2018), while the latter focuses on explaining the model’s prediction for a specific instance (Arrieta et al., 2020). The studied problem of our paper belongs to local explanations. This line of work mainly focuses on the following three directions: saliency map (Chattopadhay et al., 2018; Jalwana et al., 2021; Kim et al., 2022), concept attribution (Ghorbani et al., 2019; Kim et al., 2018; Kolek et al., 2022), and model distillation (Ge et al., 2021; Tan et al., 2018).
@@
Counterfactual Explanations As a branch of post-hoc explanations, CE has emerged as a highly influential approach in recent years (Pawelczyk et al., 2022), especially in computer vision (Jeanneret et al., 2023). Some methods generate counterfactual samples by locating key regions that distinguish two contrasting images and substituting the relevant region from one image into the other (Goyal et al., 2019; Wang et al., 2021). Other methods utilize the gradient of the input image with respect to the target label while also leveraging robust models to generate human-recognizable modifications (Schut et al., 2021; Zhu et al., 2021). Additional approaches, including prototype-based algorithms (Van Looveren & Klaise, 1907) and invertible convolutional networks (Hvilshøj et al., 2021), have also shown promising abilities in generating counterfactual examples. Generative techniques are inherently well-suited for CE. With their rapid development, an increasing number of related works have emerged, which can be broadly categorized into conditional (Singla et al., 2019; Van Looveren et al., 2021) and unconditional approaches (Khorram & Fuxin, 2022; Rodriguez et al., 2021). Diffusion models (DDPM) are cutting-edge generative models. Jeanneret et al. (2022) and (Jeanneret et al., 2023) both utilize DDPM to generate counterfactual images. The difference lies in that the former modifies the sampling process of the DDPM, while the latter employs DDPM as a regularizer.
@@
Adversarial Attacks Adversarial attacks and CEs share a similar objective: flipping the model’s prediction, which has led some studies to focus on exploring the relationship between them (Akhtar et al., 2021; Ignatiev et al., 2019; Pawelczyk et al., 2022). However, the key difference lies in the fact that adversarial attacks on undefended models often result in imperceptible noise rather than semantically meaningful changes. Adversarial attacks include white-box and black-box settings. In white-box attacks, the attacker has access to the model’s architecture and data (Costa et al., 2024). It typically generates adversarial examples by utilizing the gradient of the loss function with respect to the input (Goodfellow et al., 2014; Madry et al., 2017; Moosavi-Dezfooli et al., 2016). In contrast, black-box attacks (Andriushchenko et al., 2020; Wicker et al., 2018) restrict the attacker to accessing only samples from an oracle or pairs of inputs and corresponding outputs.
@@
Adversarial Defense The development of adversarial attacks has, in turn, stimulated significant progress in adversarial defense. Recent works can be primarily divided into two categories: certified defense and empirical defense. Certified defense (Raghunathan et al., 2018; Cohen et al., 2019; Singla & Feizi, 2020) offer provable robustness by ensuring that the model’s predictions remain stable under perturbations constrained within a specified norm ball. However, empirical defense (Goodfellow et al., 2014; Madry et al., 2017; Zhang et al., 2019), particularly adversarial training (Madry et al., 2017), is still highly effective. It typically generates adversarial samples by using an adversarial attack first and then incorporates such samples into the training process. It is worth noting that recent work has begun to integrate causality into adversarial defense strategies (Zhang et al., 2020; Carlini & Wagner, 2017; Mitrovic et al., 2020; Zhang et al., 2021), sharing a common objective with our work.
@@ 3 Preliminaries
Causal View on Image Generation Figure 1 is a well-recognized causal graph $\mathcal{G}$ (Zhang et al., 2021) that can represent the generation processes of both the natural and counterfactual images. As shown, $C$ denotes the causal factors that directly determine the label $Y$ and predicted label $\hat{Y}$. $S$ denotes the spurious factors, which are independent of $C$, as reflected in the $v$-structure $C\rightarrow{X}\leftarrow{S}$. $X$ is the natural image, and $\Delta$ represents the semantically meaningful perturbation; together, they form the counterfactual image $X^{\prime}$. $\theta$ is the set of parameters of the DNN. It is evident from the graph that the perturbation $\Delta$ is determined by $\hat{Y}$, $X$ and $\theta$. The inclusion of $\theta$ is due to the use of white-box attacks, such as projected gradient descent (PGD). Note that $S$ can also have an influence on $Y$ because of the spurious paths in $\mathcal{G}$. When $X$ is given, there is a spurious path ${Y}\leftarrow{C}\rightarrow{X}\leftarrow{S}$ associate $Y$ and $S$.
Denoising Diffusion Probabilistic Models DDPM Ho et al. (2020) can filter out the noise introduced by traditional adversarial attacks without changing the classifier’s weights. It relies on two Markov chains, corresponding to the forward and reverse processes, respectively. The forward process gradually adds Gaussian noise to the data at the state of $t$ to generate data at the next state $t+1$. In practice, the data at time step $t$, denoted as $x_{t}$ can be directly obtained from the clean data $x_{0}$ through
@@ 4 Methodology
We propose the CECAS framework, which consists of two key components, as shown in Figure 2. The first component is the initial generation of CE images, achieved through a novel causally guided adversarial method. The second component is the causally guided refinement of the generated counterfactuals, where we adopt an image inpainting approach to better preserve the details of the unmodified regions of the original image. We detail them in the following sections.
@@ 4.1 Initial Causally-Guided Counterfactual Visual Explanation Generation
The first and central component of the proposed framework is a causally guided adversarial method to generate counterfactual visual explanations. As previously mentioned, adversarial attacks and CE share a common goal: altering a model’s prediction through minimal input perturbations. Inspired by prior work bridging the two domains (Pawelczyk et al., 2022; Jeanneret et al., 2022; Augustin et al., 2022), we adopt adversarial strategies—such as PGD attacks—to generate image-based counterfactuals. The procedure is outlined as follows:
@@
However, relying solely on the above adversarial attack-based method is often overly idealized, as these methods typically impose no constraints on spurious factors. As discussed in the Introduction, although the explained model may rely on such factors during prediction, perturbing them often does not yield meaningful or actionable explanations. Furthermore, altering these spurious factors can introduce artifacts or noise, which severely compromise the core goals of CE of providing realistic and constructive explanations.
To address this issue, it is essential to constrain changes to spurious factors. Given the central role of gradients in the above CE generation process, we aim to identify a new gradient direction—one that enables effective modification of the image while preserving spurious factors unchanged. Since our goal is to explain the classifier $f$, we are not allowed to make any modifications to it. Considering the architecture of the explained model $f$, a DNN commonly viewed—following (Bengio et al., 2013)— as a combination of a feature extraction module and a linear classifier module, we learn to capture the spurious factors out of the representations $R$ learned by the feature extraction module (details of how to capture $S$ from $R$ are given in the final part of this subsection). With the captured $S$, we use the original data to train an auxiliary classifier $g$ which predicts the label $Y$ based on the spurious factor. More specifically, after training, $g$ can make predictions as follows: $\hat{z}=g(s),\hat{z}^{\prime}=g(s^{\prime})$, where $\hat{z}$ denotes the logit predicted by $g$ from the spurious factor $s$ in the original image, while $\hat{z}^{\prime}$ is the logit predicted by $g$ from the spurious factor $s^{\prime}$ in the modified image. We can consider $z$ to capture the spurious factors that are most relevant to the label YY—that is, the spurious features most likely to be unintentionally altered during counterfactual generation. On top of this, we introduce a penalty loss $L_{S}$ that encourages these spurious factors to remain invariant during the counterfactual generation process, and thereby we compute the gradient of the penalty loss with respect to the image and incorporate its direction into the guidance for image modification:
@@
where $L_{S}$ can typically be implemented as the soft cross-entropy loss, which takes two logits as input. The purpose of this loss is to align the spurious distributions between the natural and counterfactual representations by penalizing the deviation of spurious factors in the counterfactuals from those in the original inputs. In summary, the generation of counterfactual images involves modifying the images via adversarial attack, while preserving the spurious factors through a causal perspective.
Furthermore, inspired by the effectiveness of diffusion models in enhancing CE image quality (Jeanneret et al., 2022; Augustin et al., 2022), we integrate diffusion models to filter out high-frequency, semantically meaningless noise introduced by adversarial attacks. Specifically, we apply the forward process of DDPM to add Gaussian noise to the image first, and then use the reverse process to denoise it before feeding it into the classifier for loss computation. In this work, we define the joint classifier $F$ as the combination of DDPM and the base classifier $f$. The generation of counterfactual images can be simplified and reformulated in an optimization form:
@@
The only remaining problem is how to obtain the spurious factor $S$. Fortunately, numerous recent works have explored disentangling causal factors from spurious ones (Lv et al., 2022; Creager et al., 2021). Building on the causal model illustrated in Figure 1, we adopt the widely used assumption that the causal factor $C$ is independent of the spurious factor $S$. In this work, similar to (O’Shaughnessy et al., 2020), we pre-train a VAE $h$ to disentangle $C$ and $S$ from the latent space. $h$ takes the representation $R$ as input and is optimized with two objectives: (1) maximizing the causal influence of $C$ on $Y$; and (2) ensuring that $C$ and $S$, when jointly passed through the decoder, can faithfully reconstruct the distribution of $R$, thereby decomposing the latent space into two independent components.
4.2 Causally Post Refinement